OpenStudy (anonymous):

3/4x-1/2=-2/5x+1=? show work

OpenStudy (mayliz):

x = 23/30

OpenStudy (anonymous):

did you multiply all fractions by20/1

OpenStudy (anonymous):

make the equation 'equal': \[15/20x - 1/2 = -18/20x + 1\] write everything to one side: \[23/20x-3/2=0\] solve it: \[23/20x=3/2\] \[x=3/2 \times 20/23 = 60/46 = 30/23\]